OpenStudy (anonymous):

*MEDAL* can someone check this? There are 40 volunteers for the research study on the Power Pill. Each subgroup of the study will contain 10 participants. Determine how many ways these participants can be selected and explain your method. We will use a combination here since the order does NOT matter; we simply want to arrange the participants into groups of 10. 40c10= 40!/10!(40-10)! 40!/10!30!= 3075990524006400/3628800 Answer- 847660528

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

question is a bit unclear is it one subgroup of ten or four subgroups of ten?

OpenStudy (misty1212):

if it is one subgroup of ten then you are correct \[_{40}C_{10}\]

OpenStudy (misty1212):

\[\frac{40\times 39\times 38\times 37\times 36\times 35\times 34\times 33\times 32\times 31}{10\times 9\times 8\times 7\times 6\times 5\times 4\times 3\times 2}\] cancel first multiply last